OpenStudy (anonymous):

different one its solve for a: 6a+5/3 - 2a+2/2 =8/2

OpenStudy (campbell_st):

if the question is \[\frac{6a + 5}{3} - \frac{2a + 2}{2} = \frac{8}{2}\] multiply every term by 6 to eliminate the fractions \[2(6a + 5) - 3(2a +2) = 3 \times8\] expand and collect like terms 12a + 10 - 6a - 6 = 24 6a + 4 = 24 solve for a